OpenStudy (vortish):

the circumference of a circle is 16pi find the radius and area

OpenStudy (anonymous):

circumference=2*pi* radius thus radius =circumference /(2*pi*) =16pi /2pi =8 and area =pi *radius*radius =pi*8*8=64pi